Is the squence geometric? If so identify the common ratio 2, -4, -16, -36, ...
Found 2 solutions by Edwin McCravy, math_tutor2020:
Answer by Edwin McCravy(20056)


To find out if it is a geometric sequence, 
we find out whether this is completely true:







That's false, so it's not a geometric sequence.

Edwin

Answer by math_tutor2020(3817)


Answer: Not geometric

We can determine this by dividing each term by its previous term.
r = term2/term1 = -4/2 = -2
r = term3/term2 = -16/(-4) = 4
r = term4/term3 = -36/(-16) = 2.25
Technically we only need to do two calculations, but it doesn't hurt to show all three.
The results don't match up, so we don't have a geometric sequence.